A shopper pushes a grocery cart for a distance 19 M at a constant speed on level ground, against a 25 N frictional force. He pushes in a direction 25.5 degrees below the horizontal.


1
Expert's answer
2021-12-07T12:09:47-0500

(a) What is the work done on the cart by friction?


"W_f=-25\cdot19=-475\ (J)"


(b) What is the work done on the cart by the gravitational force?


"W_g=0\ (J)"


(c) What is the work done on the cart by the shopper?


"W_s=-W_f=475\ (J)"


(d) Find the forces the shopper exerts, using energy considerations.


"F\cdot\cos25.5°-25=0\to F=27.7\ (N)"


(e) what is the total work done on the cart?


"W_t=475-475+0=0\ (J)"